2. Find the total flux ΦE passing through the sides of a cube 1.00 m on a side if a point charge of 5.00 × 10−6 C is located at its center.


1
Expert's answer
2021-12-07T20:53:06-0500

Using Gauss's law "\phi=\frac{Q}{\Epsilon}"


"\phi_T=\frac{5×10^{-6}}{8.854×10^{-12}}"


"=5.647\u00d710^5Nm^2\/C"


For one side "\phi=\frac{1}{6}×5.647×10^5"

"=9.412\u00d710^4Nm^2\/C"


For all sides "=6\phi"

"=5.647\u00d710^5Nm^2\/C"